# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 3.1 Bounds in expectation
Let ˜2.2, ˜2.3, and ˜3.1. Consider the momentum stochastic gradient method (SGDM) and the extended 2D recursion (A.13). Then when the step-sizes $\gamma_{t}$ satisfies
$\gamma_{t}\leq\frac{\mu(1-\beta)^{2}}{4L^{2}},$ (3.4)
the following tracking error bound holds in expectation for SGD with momentum:
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{4}{(1-\beta)^{2}}\rho^{2(t+1)}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{48(1+\beta+\gamma L)^{2}}{(1-\beta)^{2}}\cdot\frac{\Delta^{2}}{(1-\rho)^{2}}+\frac{48}{(1-\beta)^{2}}\cdot\frac{\sigma^{2}\gamma^{2}}{1-\rho^{2}}.$ (3.5)
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}\exp\left(-\frac{\gamma\mu}{1-\beta}(t+1)\right)\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{192(2+\beta)^{2}}{\gamma^{2}\mu^{2}}\Delta^{2}+\frac{96\sigma^{2}\gamma}{\mu(1-\beta)}.$ (3.6)
Theorem˜3.3 admits the same three-way decomposition as SGD: an exponentially decaying optimization contraction plus two irreducible floors from drift and gradient noise, but with a sharper dependence on the momentum parameter. Notably, increasing momentum amplifies sensitivity to initialization through the prefactor $(1-\beta)^{-2}$ and inflates the steady-state noise-driven tracking error by $(1-\beta)^{-1}$, consistent with the interpretation of momentum as increasing the effective memory of stochastic gradients. Moreover, the exponential rate depends on $\gamma\mu/(1-\beta)$. Under standard stability choices $\gamma\lesssim 1/L$, one has $\gamma\mu\asymp 1/\kappa$ with $\kappa:=L/\mu$ so ill-conditioned objectives can exhibit a long “burn-in” before the contraction becomes negligible. This effect is particularly relevant for deep networks, whose curvature spectra are widely observed to be highly ill-conditioned with many near-flat directions (effectively small $\mu$), so the initialization term can dominate over the short horizons typical of nonstationary regimes [SBL17, JKB+19, GKX19, PAP20]. Finally, while Nesterov-style acceleration improves deterministic gradient descent from $O(\kappa)$ to $O(\sqrt{\kappa})$ iterations on smooth strongly convex problems [NES83, NES14], our bound highlights that in drifting stochastic settings, performance is governed by tracking and variance floors rather than deterministic bias decay. Consequently, the classical $\sqrt{\kappa}$ acceleration effect does not need to take effect, and aggressive momentum can primarily raise the noise floor and prolong the burn-in time. This becomes clear in light of Theorem˜3.4 which illustrates that momentum induces much longer burn-in periods:
Figure 1: Trade-off between drift tracking and stochastic noise under constant stepsize (expectation bounds). For SGD, the expected tracking error decomposes into an exponentially decaying initialization term plus two irreducible steady-state floors: a noise floor $\asymp\sigma^{2}\gamma/\mu$ and a drift-induced tracking floor $\asymp\Delta^{2}/(\mu^{2}\gamma^{2})$. For SGDM, momentum increases sensitivity to initialization by a factor $\asymp(1-\beta)^{-2}$ and inflates the noise floor by $\asymp(1-\beta)^{-1}$, while the drift floor retains the same $\gamma^{-2}$ scaling (up to constants). Moreover, stability requires $\gamma\leq\mu(1-\beta)^{2}/(4L^{2})$, so large $\beta$ can indirectly worsen drift tracking by forcing smaller admissible stepsizes. Together, these effects formalize when momentum helps in stationary regimes yet becomes fragile under drift, and when SGD is provably more robust.
@@
Theorem 3.4 (Time to reach the asymptotic tracking error in expectation for SGD with momentum).
@@
Assume ˜2.2 ˜2.3, and ˜3.1 and let $\beta\in[0,1)$. Consider SGD with momentum (SGDM). Assume a constant stepsize $\gamma_{t}\equiv\gamma$ satisfying $\gamma\leq\mu(1-\beta)^{2}/(4L^{2})$, and set

### v3

@@ C.2 Proof for SGDM tracking error bound
Let ˜2.2, ˜2.3 hold, and $\beta_{1}+\beta_{2}=\beta$ and $\beta_{1}\beta_{2}=0$ for fixed $\beta\in[0,1)$. Consider (SGDM) and the extended 2D recursion (13). Then when the step-sizes $\gamma_{t}$ satisfies
$\gamma_{t}\leq\frac{\mu(1-\beta)^{2}}{4L^{2}}$ (39)
$\begin{bmatrix}\|\widehat{\bm{\theta}}_{t+1}\|^{2}\\ \|\check{\bm{\theta}}_{t+1}\|^{2}\end{bmatrix}\leq\begin{bmatrix}a&b\\ c&d\end{bmatrix}\begin{bmatrix}\|\widehat{\bm{\theta}}_{t}\|^{2}\\ \|\check{\bm{\theta}}_{t}\|^{2}\end{bmatrix}$, (40)
where
$a=1-\frac{\gamma_{t}\mu}{1-\beta},\;\;b=\frac{\gamma_{t}{\beta^{\prime}}^{2}L^{2}}{\mu(1-\beta)},\;\;c=\frac{2\gamma_{t}^{2}L^{2}}{(1-\beta)^{3}},\;\;d=\beta+\frac{2\gamma_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{(1-\beta)^{3}}$ (41)
and the coefficient matrix is stable i.e.
$\rho\left(\begin{bmatrix}a&b\\ c&d\end{bmatrix}\right)<1.$ (42)
Proof of Corollary˜C.3.
Define the energy vector as follows:
$\mathbb{s}_{t}=\begin{bmatrix}\|\widehat{\bm{\theta}}_{t}\|^{2}\\ \|\check{\bm{\theta}}_{t}\|^{2}\end{bmatrix}\in\mathbb{R}_{\geq 0}^{2}.$ (43)
We will show the pathwise inequality $\mathbb{s}_{t+1}\leq\bm{\Gamma}_{t}\mathbb{s}_{t}$ with explicit scalar entries $a_{t},b_{t},c_{t},d_{t}$, and then get a uniform $\bm{\Gamma}$ by upper bounded $\bm{\Gamma}_{t}$ over $\forall t\geq 0$. From Lemma˜C.2, we find the first row can be written as
$\widehat{\bm{\theta}}_{t+1}=\mathbb{A}_{t}\widehat{\bm{\theta}}_{t}+\mathbb{B}_{t}\check{\bm{\theta}}_{t},$ (44)
$\|\widehat{\bm{\theta}}_{t+1}\|^{2}\leq\frac{1}{1-\tau}\|\mathbb{A}_{t}\widehat{\bm{\theta}}_{t}\|^{2}+\frac{1}{\tau}\|\mathbb{B}_{t}\check{\bm{\theta}}_{t}\|^{2}\leq\frac{\|\mathbb{A}_{t}\|^{2}}{1-\tau}\|\widehat{\bm{\theta}}_{t}\|^{2}+\frac{\|\mathbb{B}_{t}\|^{2}}{\tau}\|\check{\bm{\theta}}_{t}\|^{2}.$ (45)
$\|\widehat{\bm{\theta}}_{t+1}\|^{2}\leq\frac{(1-\eta_{t}\mu)^{2}}{1-\tau}\|\widehat{\bm{\theta}}_{t}\|^{2}+\frac{\eta_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{\tau}\|\check{\bm{\theta}}_{t}\|^{2}.$ (46)
$\|\widehat{\bm{\theta}}_{t+1}\|^{2}\leq\underbrace{(1-\eta_{t}\mu)}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;a}\|\widehat{\bm{\theta}}_{t}\|^{2}+\underbrace{\left(\eta_{t}\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}\right)}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;b}\|\check{\bm{\theta}}_{t}\|^{2}.$ (47)
We will now repeat this for the second row. We can write the second row as
$\check{\bm{\theta}}_{t+1}=-\eta_{t}\mathbb{H}_{t}\widehat{\bm{\theta}}_{t}+(\beta I_{d}+\eta_{t}\beta^{\prime}\mathbb{H}_{t})\check{\bm{\theta}}_{t}=\beta\check{\bm{\theta}}_{t}+r_{t},$ (48)
$\begin{split}\|\check{\bm{\theta}}_{t+1}\|^{2}&=\|\beta\check{\bm{\theta}}_{t}+r_{t}\|^{2}\\ &=\|\beta\check{\bm{\theta}}_{t}+(1-\beta)r_{t}/(1-\beta)^{2}\|^{2}\\ &\leq\beta\|\check{\bm{\theta}}_{t}\|^{2}+\frac{1}{1-\beta}\|r_{t}\|^{2}.\end{split}$ (49)
Now we bound $\|r_{t}\|$:
$\begin{split}\|r_{t}\|\leq\eta_{t}\|\mathbb{H}_{t}\|\|\beta^{\prime}\cdot\check{\bm{\theta}}_{t}-\widehat{\bm{\theta}}_{t}\|&\leq\eta_{t}L\left(\beta^{\prime}\|\check{\bm{\theta}}_{t}\|+\|\widehat{\bm{\theta}}_{t}\|\right).\end{split}$ (50)
Using the inequality $(a+b)^{2}\lesssim a^{2}+b^{2}$, we conclude that
$\begin{split}\|r_{t}\|^{2}\leq 2\eta_{t}^{2}L^{2}\left({\beta^{\prime}}^{2}\|\check{\bm{\theta}}_{t}\|^{2}+\|\widehat{\bm{\theta}}_{t}\|^{2}\right).\end{split}$ (51)
Therefore we find that
$\|\check{\bm{\theta}}_{t+1}\|^{2}\leq\underbrace{\left(\frac{2\eta_{t}^{2}L^{2}}{1-\beta}\right)}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;c}\|\widehat{\bm{\theta}}_{t}\|^{2}+\underbrace{\left(\beta+\frac{2\eta_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\right)}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;d}\|\check{\bm{\theta}}_{t}\|^{2}.$ (52)
Collecting (47) and (52), we find
$\begin{bmatrix}\|\widehat{\bm{\theta}}_{t+1}\|^{2}\\ \|\check{\bm{\theta}}_{t+1}\|^{2}\end{bmatrix}\leq\bm{\Gamma}_{t}\begin{bmatrix}\|\widehat{\bm{\theta}}_{t}\|^{2}\\ \|\check{\bm{\theta}}_{t}\|^{2}\end{bmatrix}$, (53)
with
$\bm{\Gamma}_{t}:=\begin{bmatrix}a&b\\ c&d\end{bmatrix}=\begin{bmatrix}1-\eta_{t}\mu&\eta_{t}\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}\\ \frac{2\eta_{t}^{2}L^{2}}{1-\beta}&\beta+\frac{2\eta_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\end{bmatrix}$. (54)
$\eta:=\sup_{t}\eta_{t}=\sup_{t}\frac{\gamma_{t}}{1-\beta}.$
Define the uniform dominating matrix
$\bm{\Gamma}:=\begin{bmatrix}1-\eta\mu&\eta\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}\\ \frac{2\eta^{2}L^{2}}{1-\beta}&\beta+\frac{2\eta^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\end{bmatrix}$. (55)
$\rho(\bm{\Gamma})\leq\max\left\{1-\eta\mu+\frac{2\eta^{2}L^{2}}{1-\beta},\beta+\eta\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}+\frac{2\eta^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\right\}.$ (56)
We can choose a step-size $\eta$ small enough to satisfy:
$\begin{cases}\frac{\eta\mu}{2}>\frac{2\eta^{2}L^{2}}{1-\beta}\\ \frac{\eta{\beta^{\prime}}^{2}L^{2}}{\mu}>\frac{2\eta^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\\ 1-\beta>\frac{2\eta{\beta^{\prime}}^{2}L^{2}}{\mu},\end{cases}$ (57)
which is equivalent to
$\eta<\min\left\{\frac{\mu(1-\beta)}{4L^{2}},\frac{1-\beta}{2\mu},\frac{\mu(1-\beta)}{{\beta^{\prime}}^{2}L^{2}}\right\}=\frac{\mu(1-\beta)}{4L^{2}}.$ (58)
$\gamma<\min\left\{\frac{\mu(1-\beta)^{2}}{4L^{2}},\frac{(1-\beta)^{2}}{2\mu},\frac{\mu(1-\beta)^{2}}{{\beta^{\prime}}^{2}L^{2}}\right\}=\frac{\mu(1-\beta)^{2}}{4L^{2}}.$ (59)
It then holds that
$\rho(\bm{\Gamma})<\max\left\{1-\frac{\gamma\mu}{2(1-\beta)},\beta+\frac{2\gamma L^{2}}{\mu(1-\beta)}\right\}\leq 1.$ (60)
In this case, $\bm{\Gamma}$ is a stable matrix. This concludes the proof. ∎
From Lemma˜C.2, we established that we have a recursive relation of the form $\mathbb{y}_{t+1}=\widetilde{\bm{B}}_{t}\mathbb{y}_{t}+\mathbb{V}^{-1}\mathbb{u}_{t}+\gamma_{t}\mathbb{V}^{-1}\mathbb{\eta}_{t+1}$. Define the original state matrices as $\bm{\Phi}(t,s):=\bm{B}_{t-1}\bm{B}_{t-2}\dots\bm{B}_{s}$ and the transformed state transition matrices as $\widetilde{\bm{\Phi}}(t,s):=\widetilde{\bm{B}}_{t-1}\widetilde{\bm{B}}_{t-2}\dots\widetilde{\bm{B}}_{s}$ for $\forall t>s$ and $\widetilde{\bm{\Phi}}(s,s):=I_{d}$. With the above established, we can convert the fact that $\bm{\Gamma}$ is stable into a bound on $\widetilde{\bm{\Phi}}$ and subsequently $\bm{\Phi}$. We will establish this below:
Corollary C.4 (Conversion from $\widetilde{\bm{\Phi}}$ to $\bm{\Phi}$ ).
Let ˜2.2, ˜2.3 hold and $\beta_{1}+\beta_{2}=\beta$ with $\beta_{1}\beta_{2}=0$ for fixed $\beta\in[0,1)$. Consider (SGDM) and the extended 22D recursion (13). Assume $\eta_{t}=\gamma_{t}/(1-\beta)$ and
$\gamma_{t}\leq\frac{\mu(1-\beta)^{2}}{4L^{2}}.$
Let $\bm{\Gamma}_{t}$ be defined by
$\bm{\Gamma}_{t}=\begin{bmatrix}a_{t}&b_{t}\\ c_{t}&d_{t}\end{bmatrix}=\begin{bmatrix}1-\eta_{t}\mu&\eta_{t}\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}\\ \frac{2\eta_{t}^{2}L^{2}}{1-\beta}&\beta+\frac{2\eta_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\end{bmatrix},\qquad\beta^{\prime}=\beta\beta_{1}+\beta_{2}.$
Define
$\bar{\rho}\;:=\;\sup_{t\geq 0}\sqrt{\|\bm{\Gamma}_{t}\|_{1}},\qquad\text{where}\quad\|\bm{\Gamma}_{t}\|_{1}=\max\{a_{t}+c_{t},\ b_{t}+d_{t}\}.$
Then $\bar{\rho}<1$ and for all $t\geq\alpha$,
$\|\bm{\Phi}(t,\alpha)\|_{\mathrm{op}}\leq\frac{4}{1-\beta}\,\bar{\rho}^{\,t-\alpha}.$
Proof of Corollary˜C.4.
First note that $\bm{y}_{t}=\widetilde{\bm{\Phi}}(t,\alpha)\bm{y}_{\alpha}$. Recall from Lemma˜C.2 that we have
$\|\bm{y}_{t}\|^{2}=\|\widehat{\bm{\theta}}_{t}\|^{2}+\|\check{\bm{\theta}}_{t}\|^{2}=\mathbf{1}^{\top}\bm{s}_{t},$ (61)
where $\bm{s}_{t}:=[\|\widehat{\bm{\theta}}_{t}\|^{2};\|\check{\bm{\theta}}_{t}\|^{2}]\in\mathbb{R}_{+}^{2}$. In Corollary˜C.3, we showed that for all $k\geq 0$,
$\bm{s}_{k+1}\leq\bm{\Gamma}_{k}\bm{s}_{k}\qquad\text{(entrywise)}.$ (62)
Iterating (62) from $\alpha$ to t1t-1 gives
$\bm{s}_{t}\leq\bm{\Gamma}_{t-1}\bm{\Gamma}_{t-2}\cdots\bm{\Gamma}_{\alpha}\,\bm{s}_{\alpha}\qquad\text{(entrywise)}.$ (63)
Taking $\mathbf{1}^{\top}(\cdot)$ and using nonnegativity yields
$\|\bm{y}_{t}\|^{2}=\mathbf{1}^{\top}\bm{s}_{t}\leq\mathbf{1}^{\top}\Big(\bm{\Gamma}_{t-1}\cdots\bm{\Gamma}_{\alpha}\Big)\bm{s}_{\alpha}\leq\|\bm{\Gamma}_{t-1}\cdots\bm{\Gamma}_{\alpha}\|_{1}\ \|\bm{s}_{\alpha}\|_{1}.$ (64)
Since $\|\bm{s}_{\alpha}\|_{1}=\mathbf{1}^{\top}\bm{s}_{\alpha}=\|\bm{y}_{\alpha}\|^{2}$, we obtain
$\|\bm{y}_{t}\|^{2}\leq\|\bm{\Gamma}_{t-1}\cdots\bm{\Gamma}_{\alpha}\|_{1}\ \|\bm{y}_{\alpha}\|^{2}.$ (65)
By submultiplicativity of the induced $1$-norm,
$\|\bm{\Gamma}_{t-1}\cdots\bm{\Gamma}_{\alpha}\|_{1}\leq\prod_{k=\alpha}^{t-1}\|\bm{\Gamma}_{k}\|_{1}\leq\Big(\sup_{k\geq 0}\|\bm{\Gamma}_{k}\|_{1}\Big)^{t-\alpha}.$ (66)
Define
$\bar{\rho}\;:=\;\sqrt{\sup_{k\geq 0}\|\bm{\Gamma}_{k}\|_{1}}.$ (67)
Combining (65) and (67) yields
$\|\bm{y}_{t}\|\leq\bar{\rho}^{\,t-\alpha}\|\bm{y}_{\alpha}\|.$ (68)
Therefore,
$\|\widetilde{\bm{\Phi}}(t,\alpha)\|_{\mathrm{op}}\leq\bar{\rho}^{\,t-\alpha}.$ (69)
Since we have $\bm{z}_{t}=\bm{V}\bm{y}_{t}$, it follows that
$\bm{\Phi}(t,\alpha)=\bm{B}_{t-1}\cdots\bm{B}_{\alpha}=\bm{V}\widetilde{\bm{\Phi}}(t,\alpha)\bm{V}^{-1}.$
Hence
$\|\bm{\Phi}(t,\alpha)\|_{\mathrm{op}}\leq\|\bm{V}\|_{\mathrm{op}}\|\bm{V}^{-1}\|_{\mathrm{op}}\|\widetilde{\bm{\Phi}}(t,\alpha)\|_{\mathrm{op}}\leq\|\bm{V}\|_{\mathrm{op}}\|\bm{V}^{-1}\|_{\mathrm{op}}\,\bar{\rho}^{\,t-\alpha}.$ (70)
From (9), one can show that $\|\bm{V}[x;y]\|\leq\sqrt{2}(\|x\|+\|y\|)\leq 2\|[x;y]\|$, hence $\|\bm{V}\|_{\mathrm{op}}\leq 2$. This also implies $\|\bm{V}^{-1}\|_{\mathrm{op}}\leq 2/(1-\beta)$. Plugging into (70) yields
$\|\bm{\Phi}(t,\alpha)\|_{\mathrm{op}}\leq\frac{4}{1-\beta}\,\bar{\rho}^{\,t-\alpha}.$ (71)
It remains to show that $\bar{\rho}<1$. Note that
$\|\bm{\Gamma}_{t}\|_{1}=\max\{a_{t}+c_{t},\ b_{t}+d_{t}\},$
where
$a_{t}+c_{t}=1-\eta_{t}\mu+\frac{2\eta_{t}^{2}L^{2}}{1-\beta}\leq 1-\frac{\eta_{t}\mu}{2},\qquad\text{since }\ \frac{2\eta_{t}^{2}L^{2}}{1-\beta}\leq\frac{\eta_{t}\mu}{2}\ \text{ under }\ \gamma_{t}\leq\frac{\mu(1-\beta)^{2}}{4L^{2}}.$
Moreover, using ${\beta^{\prime}}^{2}\leq 1$,
$b_{t}+d_{t}=\beta+\eta_{t}\frac{{\beta^{\prime}}^{2}L^{2}}{\mu}+\frac{2\eta_{t}^{2}{\beta^{\prime}}^{2}L^{2}}{1-\beta}\leq\beta+\eta_{t}\frac{L^{2}}{\mu}+\frac{2\eta_{t}^{2}L^{2}}{1-\beta}.$
With $\eta_{t}=\gamma_{t}/(1-\beta)$ and $\gamma_{t}\leq\mu(1-\beta)^{2}/(4L^{2})$, we have
$\eta_{t}\frac{L^{2}}{\mu}\leq\frac{1-\beta}{4},\qquad\frac{2\eta_{t}^{2}L^{2}}{1-\beta}\leq\frac{1-\beta}{8},$
and therefore
$b_{t}+d_{t}\leq\beta+\frac{3}{8}(1-\beta)=\frac{3}{8}+\frac{5}{8}\beta<1.$
Consequently $\sup_{t\geq 0}\|\bm{\Gamma}_{t}\|_{1}<1$, and thus $\bar{\rho}<1$ by (67). This completes the proof. ∎
In light of these results, we can finally prove an expectation bound on the tracking error of (SGDM):
Corollary C.5 (Tracking error bound in expectation for (SGDM)).
Let ˜2.2, ˜2.3, and ˜3.1. Consider (SGDM) and the extended 2D recursion (13). Then when the step-sizes $\gamma_{t}$ satisfies
$\gamma_{t}\leq\frac{\mu(1-\beta)^{2}}{4L^{2}},$ (72)
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}\rho^{2(t+1)}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{48(1+\beta+\gamma L)^{2}}{(1-\beta)^{2}}\cdot\frac{\Delta^{2}}{(1-\rho)^{2}}+\frac{48}{(1-\beta)^{2}}\cdot\frac{\sigma^{2}\gamma^{2}}{1-\rho^{2}}.$ (73)
@@
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}\exp\left(-\frac{\gamma\mu}{1-\beta}(t+1)\right)\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{192(2+\beta)^{2}}{\gamma^{2}\mu^{2}}\Delta^{2}+\frac{96\sigma^{2}\gamma}{\mu(1-\beta)}.$ (74)
Proof of Corollary˜C.5.
For the analysis, we will denote $C_{\beta}:=4/(1-\beta)$. From Lemma˜C.2, we showed that (SGDM) follows a 2D recursive system as follows:
$\underbrace{\begin{bmatrix}\widetilde{\bm{\theta}}_{t+1}\\ \widetilde{\bm{\theta}}_{t}\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{z}_{t+1}}=\underbrace{\begin{bmatrix}\mathbb{J}_{t}&\mathbb{K}_{t}\\ I_{d}&0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{B}_{t}}\underbrace{\begin{bmatrix}\widetilde{\bm{\theta}}_{t}\\ \widetilde{\bm{\theta}}_{t-1}\end{bmatrix}}_{{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{z}_{t}}}+\underbrace{\begin{bmatrix}\mathbb{b}_{t}\\ 0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\mathbb{u}_{t}}+\gamma_{t}\underbrace{\begin{bmatrix}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ 0\end{bmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\bm{\eta}_{t+1}}$ (75)
where
$\begin{split}&\mathbb{J}_{t}=(1+\beta)I_{d}-\gamma_{t}(1+\beta_{1})\mathbb{H}_{t}\\ &\mathbb{K}_{t}=-\beta I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t}\\ &\mathbb{b}_{t}=-(I_{d}-\gamma_{t}\mathbb{}H_{t})\Delta_{t}-\mathbb{K}_{t}\Delta_{t-1}.\end{split}$ (76)
Iterating (75) gives us the identity:
$\mathbb{z}_{t+1}=\bm{\Phi}(t+1,0)\mathbb{z}_{0}+\sum_{k=0}^{t}\bm{\Phi}(t+1,k+1)\mathbb{u}_{k}+\sum_{k=0}^{t}\bm{\Phi}(t+1,k+1)\gamma_{k}\mathbb{\eta}_{k+1}.$ (77)
Define the following quantities:
$A_{t+1}:=\bm{\Phi}(t+1,0)\mathbb{z}_{0},\;\;D_{t+1}:=\sum_{k=0}^{t}\bm{\Phi}(t+1,k+1)\mathbb{u}_{k},\;\;N_{t+1}:=\sum_{k=0}^{t}\bm{\Phi}(t+1,k+1)\gamma_{k}\bm{\eta}_{k+1}.$ (78)
Then we have that
$\mathbb{z}_{t+1}=A_{t+1}+D_{t+1}+N_{t+1}.$ (79)
$\mathbb{E}\|\mathbb{z}_{t+1}\|^{2}\leq 3\mathbb{E}\|A_{t+1}\|^{2}+3\mathbb{E}\|D_{t+1}\|^{2}+3\mathbb{E}\|N_{t+1}\|^{2}.$ (80)
We will now proceed with bounding each of these quantities.
Bounding $\mathbb{E}\|A_{t+1}\|^{2}$:
From the stability established in Corollary˜C.4, we have
$\|A_{t+1}\|=\|\bm{\Phi}(t+1,0)\mathbb{z}_{0}\|\leq C_{\beta}\rho^{t+1}\|\mathbb{z}_{0}\|.$ (81)
Bounding $\mathbb{E}\|D_{t+1}\|^{2}$:
First simply note that $\|\mathbb{u}_{t}\|=\|\mathbb{b}_{t}\|$. Thus it suffices to bound $\|\mathbb{b}_{t}\|$. By definition, we have that $\mathbb{b}_{t}=-(I_{d}-\gamma_{t}\mathbb{}H_{t})\Delta_{t}-\mathbb{K}_{t}\Delta_{t-1}$. Since $\mu I_{d}\preceq\mathbb{H}_{t}\preceq LI_{d}$ and $\gamma_{t}\leq 1/L$, the eigenvalues of $I_{d}-\gamma_{t}\mathbb{H}_{t}$ lie in $1-\gamma_{t}L,1-\gamma_{t}\mu]\subset[0,1]$ so we have that $\|I_{d}-\gamma_{t}\mathbb{H}_{t}\|\leq 1$. Also note that since $\mathbb{K}_{t}=-\beta I_{d}+\gamma_{t}\beta_{1}\mathbb{H}_{t}$, we have $\|\mathbb{K}_{t}\|\leq\beta+\gamma_{t}L$. Thus putting everything together, we find
$\|\mathbb{b}_{t}\|\leq\|I_{d}-\gamma_{t}\mathbb{H}_{t}\|\|\Delta_{t}\|+\|K_{t}\|\|\Delta_{t-1}\|\leq(1+\beta+\gamma_{t}L)\Delta.$ (82)
Using triangle inequality and the bound on $\|\mathbb{b}_{t}\|$, we have
$\|D_{t+1}\|^{2}\leq\sum_{k=0}^{t}\|\bm{\Phi}(t+1,k+1)\|^{2}\|\mathbb{b}_{k}\|^{2}\leq C_{\beta}^{2}\Delta^{2}(1+\beta+\gamma_{t}L)^{2}\left(\sum_{k=0}^{t}\rho^{t-k}\right)^{2}.$ (83)
$\mathbb{E}\|D_{t+1}\|^{2}\leq C_{\beta}^{2}(1+\beta+\gamma_{t}L)^{2}\frac{\Delta^{2}}{(1-\rho)^{2}}.$ (84)
Bounding $\mathbb{E}\|N_{t+1}\|^{2}$:
We will work with the recursion that $N_{t+1}$ satisfies:
$N_{t+1}=\mathbb{B}_{t}N_{t}+\gamma_{t}\bm{\eta}_{t+1}$ (85)
with $N_{0}=0$. Multiplying by $\mathbb{V}^{-1}$, we find that
$n_{t+1}=\widetilde{\mathbb{B}}_{t}n_{t}+\gamma_{t}\widetilde{\bm{\eta}}_{t+1}$ (86)
$\mathbb{E}[\|n_{t+1}\|^{2}]=\|\widetilde{\mathbb{B}}_{t}n_{t}\|^{2}+\gamma_{t}^{2}\mathbb{E}[\|\widetilde{\bm{\eta}}_{t+1}\|^{2}].$ (87)
From the stability established in Corollary˜C.3, we have that $\|\widetilde{\mathbb{B}}_{t}n_{t}\|^{2}\leq\rho\|n_{t}\|^{2}$. Also note that from Lemma˜C.2, we have an explicit form for $\widetilde{\bm{\eta}}_{t+1}$:
$\widetilde{\bm{\eta}}_{t+1}=\frac{1}{1-\beta}\begin{bmatrix}\bm{\xi}_{t+1}(\bm{\psi}_{t})\\ \bm{\xi}_{t+1}(\bm{\psi}_{t})\end{bmatrix}$. (88)
Under the boundedness assumption (˜C.1), we have that
$\mathbb{E}[\|\widetilde{\bm{\eta}}_{t+1}\|^{2}]\leq\frac{2\sigma^{2}}{(1-\beta)^{2}}.$ (89)
Thus we find that
$\mathbb{E}[\|n_{t+1}\|^{2}]=\rho\|n_{t}\|^{2}+\frac{2\gamma_{t}^{2}\sigma^{2}}{(1-\beta)^{2}}.$ (90)
Unrolling this expression and using the fact that $N_{0}=0$, we find that
$\mathbb{E}[\|n_{t+1}\|^{2}]\leq\frac{2\gamma_{t}^{2}\sigma^{2}}{(1-\beta)^{2}(1-\rho)^{2}}.$ (91)
Thus we can conclude with:
$\mathbb{E}\|N_{t+1}\|^{2}=\|\mathbb{V}\|^{2}\mathbb{E}[\|n_{t+1}\|^{2}]\leq\frac{8\gamma_{t}^{2}\sigma^{2}}{(1-\beta)^{2}(1-\rho)^{2}}.$ (92)
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}\rho^{2(t+1)}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{48(1+\beta+\gamma L)^{2}}{(1-\beta)^{2}}\cdot\frac{\Delta^{2}}{(1-\rho)^{2}}+\frac{24}{(1-\beta)^{2}}\cdot\frac{\sigma^{2}\gamma^{2}}{1-\rho^{2}}.$ (93)
From Corollary˜C.3, we can take $\rho=1-\gamma\mu/2(1-\beta)$ and use the inequality $(1-a)^{2(t+1)}\leq e^{-2a(t+1)}$ to conclude:
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\leq\frac{48}{(1-\beta)^{2}}\exp\left(-\frac{\gamma\mu}{1-\beta}(t+1)\right)\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{192(2+\beta)^{2}}{\gamma^{2}\mu^{2}}\Delta^{2}+\frac{96\sigma^{2}\gamma}{\mu(1-\beta)}.$ (94)
Using Corollary˜C.5, we can similarly obtain an algorithmic guarantee for (SGDM):
@@
Corollary C.6 (Time to reach the asymptotic tracking error in expectation for (SGDM)).
@@
Assume ˜2.2 ˜2.3, and ˜3.1 and let $\beta\in[0,1)$. Consider (SGDM). Assume a constant stepsize $\gamma_{t}\equiv\gamma$ satisfying $\gamma\leq\mu(1-\beta)^{2}/(4L^{2})$, and set